# KompeteAI: Accelerated Autonomous Multi-Agent System for End-to-End Pipeline Generation for Machine Learning Problems

### v1

@@ Appendix D Metrics
Competition RD-agent AIDE KompeteAI cmi-detect-behavior-with-sensor-data None None None eedi-mining-misconceptions-in-mathematics 22 None 30 equity-post-HCT-survival-predictions None None None learning-agency-lab-automated-essay-scoring-2 12 19 13 llm-prompt-recovery 12 9 None lmsys-chatbot-arena None None None make-data-count-finding-data-references 11 None 18 neurips-open-polymer-prediction-2025 13 None 15 pii-detection-removal-from-educational-data None None 29 um-game-playing-strength-of-mcts-variants None None None wsdm-cup-multilingual-chatbot-arena 14 None 19
@@
Results of AutoML system on Kompete-bench are presented in the Table 9 and Table 10. We report results using the ”percent humans beaten” metric, which reflects the percentage of Kaggle leaderboard participants outperformed by each system. For every competition, each system was evaluated over three independent runs; the final score is the arithmetic mean of these runs, rounded to the nearest integer. If a system failed to generate any valid solution across all three attempts, its score is reported as None (equivalent to 0 when averaging across competitions). For individual runs where no valid submission was produced, a score of 0 was assigned. Notably, the primary factor influencing overall performance was the proportion of valid submissions: systems that consistently generated correct code achieved substantially higher scores.

### v2

@@ Appendix E Metrics
Competition RD-agent AIDE ML-Master MLE-STAR KompeteAI cmi-detect-behavior-with-sensor-data None None None None None eedi-mining-misconceptions-in-mathematics 19 None 18 9 30 equity-post-HCT-survival-predictions None None None None None learning-agency-lab-automated-essay-scoring-2 9 23 15 17 13 llm-prompt-recovery 11 11 17 None None lmsys-chatbot-arena None None 14 17 15 make-data-count-finding-data-references 13 None 17 21 18 neurips-open-polymer-prediction-2025 9 None 10 11 21 pii-detection-removal-from-educational-data None None 19 13 38 um-game-playing-strength-of-mcts-variants None None None None None wsdm-cup-multilingual-chatbot-arena 14 None 8 10 19
@@
Results of the AutoML system on Kompete-bench are presented in Table 9 and Table 10. We report results using the ”percent humans beaten” metric, which reflects the percentage of Kaggle leaderboard participants outperformed by each system. For every competition, each system was evaluated over three independent runs; the final score is the arithmetic mean of these runs, rounded to the nearest integer. If a system failed to generate any valid solution across all three attempts, its score is reported as None (equivalent to 0 when averaging across competitions). For individual runs where no valid submission was produced, a score of 0 was assigned. Notably, the primary factor influencing overall performance was the proportion of valid submissions: systems that consistently generated correct code achieved substantially higher scores.
Appendix F Analysis
F.1 Performance Assessment of the Scoring Model
Figure A4: Comparison between predicted scores and actual validation scores after min–max normalization within each competition. Each point corresponds to a solution from a distinct competition. The solid black diagonal represents the ideal case where predicted and actual scores are perfectly aligned.
To evaluate the proposed scoring model, we conducted experiments across multiple benchmark competitions. Figure A4 compares predicted performance estimates with actual scaled scores obtained from validation splits of the data, covering tasks from image recognition to NLP and structured data. To account for the fact that competitions differ substantially in their metric ranges, we normalize scores within each competition using min–max scaling:
$s^{\text{scaled}}_{i}=\frac{s^{\text{val}}_{i}-\min_{j}s^{\text{val}}_{j}}{\max_{j}s^{\text{val}}_{j}-\min_{j}s^{\text{val}}_{j}},$
where $s^{\text{val}}_{i}$ denotes the validation score of solution $i$ within a competition. This adjustment ensures that relative differences are preserved while eliminating distortions caused by heterogeneous metric ranges, which would otherwise make cross-task comparisons less interpretable.
The results show a clear positive correlation between predicted and real scores. The model not only distinguishes strong solutions from weaker ones but also preserves relative ordering across a wide range of performance levels, which is critical for ranking-based exploration. At the same time, deviations from the diagonal indicate systematic tendencies: predictions for top-performing submissions are slightly conservative, while weaker candidates are occasionally overestimated. This asymmetry suggests that the model captures the general performance landscape but remains imperfect in calibrating extreme cases. Importantly, such biases are less problematic for the intended use case, where relative ranking — rather than precise absolute accuracy — guides the search process.
Two caveats remain: prediction errors may accumulate during long-running searches, and the evaluation is restricted to Kaggle-style benchmarks. Extending validation to real scientific discovery tasks represents an important future direction.
F.2 Impact of Competition and Literature Retrieval in RAG
Configuration Variant Relative Performance (%)
Without competition knowledge 86.7
Without arXiv knowledge 92.2
Without any RAG input 83.4
Table 12: Ablation of retrieval-augmented knowledge sources on a subset of benchmark competitions. Reported values indicate the relative performance by competition metric of each ablated configuration compared to the complete RAG-enabled system (treated as the 100% reference point).
The ablation analysis in Table 12 shows how different retrieval sources contribute to overall system performance. Removing competition data reduces relative accuracy to 86.7%, suggesting that competition write-ups often contain concrete heuristics and task-specific tricks that transfer effectively to our benchmark problems.
Excluding arXiv papers has a milder effect (92.2%). This is consistent with the idea that publications tend to capture broader methodological advances. While such knowledge is not always directly applicable to competition-style tasks, it still provides useful algorithmic patterns that enhance performance.
The largest drop is observed when retrieval augmentation is disabled entirely (83.4%). This indicates that external knowledge, regardless of source, is a key component of the approach. Together, competition reports and academic papers complement each other: the former supply practical, domain-oriented techniques, while the latter contribute generalizable methods and perspectives.
F.3 Distribution of Successfully Integrated Research Ideas
Figure A5: Average outcomes of the debugging process. The chart illustrates the proportion of successfully integrated ideas, categorized by their target AutoML stage — Model Training (MT) or Feature Engineering (FE) — versus the proportion of ideas that were ultimately discarded as ”Failed”.
Figure A5 summarizes the outcomes of the debugging and integration pipeline. Overall, 80.4% of proposed ideas were successfully incorporated into the research tree, confirming the reliability of debugging mechanisms. Only 19.6% of ideas failed to pass validation, that there are significantly fewer ideas that our system successfully debug.